# DashCLIP: Leveraging multimodal models for generating semantic embeddings for DoorDash

### v1

@@ Abstract.
Product Embeddings, Vision-Language Pre-training, Click-through Rate Prediction
@@ 1. Introduction
Personalized recommendation (Cheng and Cantú-Paz, 2010; Pancha et al., 2022; Zhang et al., 2024; Baltescu et al., 2022), ubiquitous in e-commerce and social network platforms on the internet, is to find out relevant entities given the user preferences and context-dependent needs. This aligns with the business objective of creating user value and fostering long-term customer loyalty. Deep learning models have emerged as the foundation of modern machine learning systems in the space, excelling at capturing complex relationships between user preferences, product attributes, and contextual signals (Cheng et al., 2016; Guo et al., 2017; Wang et al., 2017, 2021). A key strength of deep learning models is their ability to learn embedding representations for categorical features while also leveraging pre-trained embeddings generated by more sophisticated models that are infeasible to serve online at web scale. In practice, deploying embedding-based approaches across various e-commerce applications presents several challenges: (1) Embeddings generated from historical user interactions may fail to capture the rich semantic structure of product attributes, resulting in biased exposure and diminished diversity. (2) Pretrained embeddings may struggle to generalize effectively to out-of-distribution scenarios, restricting their applicability across diverse product and user tasks in e-commerce applications.
@@
In this work, we present an embedding generation and alignment framework that we develop within the context of DoorDash’s Consumer Packaged Goods (CPG) business which encompasses the delivery of groceries, retail products, alcohol, electronics, pharmaceuticals, and more. Our framework, named as DashCLIP, is built on top of the recent methodology of pre-training multi-modal models using contrastive learning (Radford et al., 2021; Li et al., 2022). DashCLIP leverages these key design choices to create semantic entity representations which are generalizable to different functional requirements:
@@
Multi-Modality Encodings: In e-commerce, products are typically associated with both textual and visual information. Instead of processing these modalities separately, it is essential to integrate the rich, yet unstructured, data from both sources to create a unified representation. In this work, we leverage contrastive learning on DoorDash’s product catalog to approximate human-like understanding of products, capturing the complementary information from each modality to improve recommendation performance.
@@
Domain Adaptation: Transfer learning provides a solution by enabling fine-tuning of the off-the-shelf models for domain-specific applications, as demonstrated in prior work on fine-tuned image embeddings (Zhai et al., 2019) and fine-tuned text embeddings (Zhuang et al., 2020). To achieve this, we perform continual pre-training on a multi-modal transformer-based model using contrastive learning on DoorDash’s internal catalog data, leveraging product images and titles as inputs to adapt to DoorDash’s specific catalog distribution.
@@
Embedding Alignment: To ensure seamless integration with downstream applications, the architecture must be adaptable to accommodate new task objectives. For search recommendation use cases, we introduce a second stage of alignment by incorporating a dedicated query encoder. This encoder is designed to generate query embeddings that are learned in the same space as product embeddings, enabling more effective retrieval and ranking. This alignment strategy serves as a foundational fine-tuning approach for any general e-commerce or social media search applications.
@@
Relevance Dataset Curation: For a product search application, product-query relevance is crucial to match the user intents and product offerings. Most prior systems resolve this connection through historical user engagements like clicks or conversions, which are prone to position and selection bias. To address this, we curate a high-quality relevance dataset that combines internal human annotations with knowledge from large language models, providing robust supervision for aligning product and query embeddings.
We have successfully deployed DashCLIP to drive ads recommendation on DoorDash’s CPG surfaces. In the online A/B experiments, we observed significant gains in user engagements and other business metrics. Standard offline model metrics, such as ROC-AUC and normalized log-loss demonstrated that the design choices rewarded exceptionally well. Following this, we evaluated the generalizability of DashCLIP through aisle category prediction and product-query relevance prediction which achieved significant gains over pre-trained vision-language models.
@@
The remainder of this paper is structured as follows: Section 2 covers the data collection and curation process; Section 3 details the architecture of our multi-modal encoders, along with the training procedure; Section 4 offers a thorough evaluation of our model’s performance in search ranking; Section 5 discusses potential e-commerce applications; and Section 6 & 7 concludes and reviews related work.
@@ 2.1. Catalog Dataset
The product catalog contains various types of information related to products available on DoorDash. This includes merchant-provided and internally collected data about the product. We curate a list of around 400k products and use their catalog data for our continual pre-training and evaluation tasks. While the catalog contains a lot of different fields, we find the following to be important due to their semantic value: 1) Title: The ”name” of the product as shown to the user. 2) Image: A standardized image of the product. 3) Description: Usually, a single sentence elaborating some additional information, e.g. dimensions for packaged goods. 4) Detail: Usually, a few lines or a paragraph explaining the product features which are more verbose compared to the description. 5) Aisle Category: The category of the product classified with internal taxonomy, e.g. drinks, snacks.
For the purpose of this paper, we only utilize the product title and product image to train the embeddings. We make this decision for two main reasons. First, the title and image are likely to receive the largest user attention on the platform as only these two fields are visible to the users in the category or search results page. Additionally, we incorporate images along with text, as we have observed that user engagement with products is strongly impacted by the images shown alongside the product titles. Next, while the description and detail sometimes contain additional information, we observed some data quality issues with these two fields. The detail is often framed like an advertisement for the product which makes it quite noisy to use without further processing. Additionally, from an exploratory data analysis conducted on selected 13k products, we found that the description and detail to be missing for 32% and 20% of the products respectively, which raised data coverage concerns. Lastly, we use aisle category as labels as part of the sampling strategy for contrastive learning and the additional downstream task described in Section 5.1.
@@ 2.2. Query Product Relevance Dataset
You are an AI online grocery shopping expert. Given a query that
a user searched for, and an item shown to the user, your job is to
understand the relevance of the item to the search query. We
classify the relevance of the item to the search query with
three distinct labels, that have the following definition:
Highly Relevant (relevance: 2): The item is exactly what you
would expect in search results. The item is clearly helpful to
show. This item fulfills the primary intent for the search term.
Moderately Relevant (relevance: 1): The item is a reasonable
substitute if the ideal item is not available. The item is similar to
the ideal item and fulfills the same general intent. These are often
items under a shared category but differ on specific attributes
such as brand (soda: Coke vs Pepsi), flavor (ice cream: chocolate
vs vanilla), ingredient (flour: almond flour vs coconut flour),
nutritional content (milk: 2% vs whole) and size (Drinking Water
(1 gal) vs Spring Water (16.9 oz x 35 ct)). This also includes items
under different/sister categories that fulfill the same intent, such
as toilet paper vs wet wipes (used in the bathroom) and waffles
vs pancakes (similar breakfast dishes). The item might also be
somewhat relevant if it fulfills a secondary intent for the search
term. For instance, some minority of users would have
reasonably expected a different ideal item. In numerous
instances, there will be a partial overlap between a segment
of the search query and the items name.
Irrelevant (relevance: 0): Item does not belong in the search
results for this query. The item is clearly not helpful to show
to a user.
Examples:
search query: arugula
item name: BrightFresh Micro Arugula (1.75 oz)
relevance: 2
search query: arugula
item name: Spinach Bunch (bunch)
relevance: 1
Listing 1: Prepended prompt for fine-tuned GPT model to infer query product relevance
In order to align the query embedding and product embedding in the same space, we require a relevance dataset that assigns a relevance label from {0: irrelevant, 1: moderately relevant, 2: highly relevant} to each ¡query, product¿ pair. We refer to these labels as IR, MR, and HR. We used a hybrid approach to utilize both human annotation and LLM to create such a dataset. Specifically, we have 700k human-annotated relevance labels, which is used to fine-tune a language model (GPT 3.5 as the final choice) to grade the relevance of any new ¡query, product¿ pair. In total, there are 32 million LLM-generated relevance labels for such pairs which cover more than 99% of our search volume. The distribution of the labels is {IR: 69%, MR: 20.5%, HR: 9.9%}. This process was done in late 2023 so GPT3, 3.5, 4 were evaluated. For our specific relevance task, we found out using a more sophisticated models may not guarantee enhanced performance, but fine-tuning has a notable improvement. Thus, we fine-tuned the language model with 600k human annotations and evaluated it with the remaining 100k to select the best version for inference. An example for fine-tuning is included above: The prompt specified the role of the language model and the definition of different relevance labels. After the initial prompt, a list of ¡query, product¿ examples are appended.
@@ 3.1. Model Architecture
The goal of DashCLIP is to create generalized product representations that can be applied across various downstream applications. For instance, an image de-duplication task, where merchants upload product images to the catalog, would require an image-only encoder. Similarly, an image retrieval task would involve both text-only and image-only encoders, with their embeddings aligned in the same space. Thus, our embedding models are designed to support multiple modalities, enabling them to meet the specific requirements of different tasks.
To achieve this, we adopt the vision-language pre-training framework outlined in (Li et al., 2022), which consists of two uni-modal encoders (one for image and one for text) and a third image-grounded text encoder. These encoders work together to generate product representations. To address scenarios involving user search, such as the search ranking experiments discussed in Section 4, we incorporate an additional query encoder. This query encoder utilizes a text-only transformer to process normalized free-text search terms provided by the user. The final model architecture is illustrated in Figure 1.
@@ 3.2. Model Training
We initialize the image-text Product Encoders from a pre-trained checkpoint BLIP-14M (Li et al., 2022). The authors also make available the checkpoints which are fine-tuned on COCO (Lin et al., 2014) and Flickr30K (Young et al., 2014) datasets, however, we don’t opt for these as the fine-tuning datasets comprise of open-world and human images which are quite different from product images. After the initialization, we adopt a two-staged training methodology:
@@ 3.2.1. Continual Pre-training of Product Encoders
In this first stage, we continue the pre-training of the Product Encoders on the 400k raw product images and titles from our Catalog Dataset mentioned in Section 2.1. By exposing the encoders to the product data from DoorDash’s catalog, the encoders will adapt to the characteristics and patterns of the product domain. This optimization stage aims to minimize the contrastive loss between image and text embeddings (ITC), and the matching loss of the image-text representations (ITM). We follow the same soft label creation and negative sampling strategies from (Li et al., 2022, 2021). In order to avoid overfitting the encoders, we freeze the first 8 layers of image and text encoders and early-stop at the lowest evaluation score of the matching loss.
@@ 3.2.2. Aligning Query and Product Encoders
where $C_{i}$ is the multi-modal hidden representation of the $i$-th product, $Q_{i}^{+}$ is the positive (relevant) query for the $i$-th product, $Q_{ij}^{-}$ is the $j$-th negative query among the $N$ negative samples for the $i$-th product. We average this loss over the batch size $B$. $sim$ is the cosine similarity function, and $\tau$ is the temperature parameter.
@@
We perform these two training stages to obtain the final encoders as the inference model to generate DashCLIP embeddings for any product or query.
4. Search Ranking Experiments
The DashCLIP product and query embeddings are generalized for different downstream applications. In this section, we focus on the task of click-through rate (CTR) prediction, which infers the probability of a DoorDash user engaging with the advertisement shown on the search surface (Richardson et al., 2007). We can retrieve the product embedding based on the identifier of the advertised item and the query embedding based on the normalized search term from the user. The embeddings after the final projection layer in the encoders will be integrated as input features to the core deep neural network based ranking model. We ran offline experiments with ablation studies to investigate the best model architecture to make use of the embeddings, and deployed the best candidate in the production environment to serve real users and verify the business impact. Highlights of the evaluations are as follows:
@@ 4. Search Ranking Experiments
    DashCLIP embeddings outperform those learned through a randomly initialized embedding lookup matrix for various products during end-to-end model training in offline experiments. This highlights that utilizing the semantic information of products and queries to pre-train the embeddings provides a strong prior for downstream applications.
@@
$NW$setminus $U_{PurcHist}$ DCN baseline .7456 $\pm$ .0004 .9183 $\pm$ .0006
@@
Table 1. Search Ranking Experiment. We collect the evaluation data from one week after the training data (called $NW$). Several evaluations were done to highlight different perspectives of the performance gain. The best candidate with product + query + purchase history embeddings is bolded and its improvements over the baseline were statistically significant ($p<0.05$). Also, the candidate performed better for users with purchase history ($U_{PurcHist}$) by capturing a semantic profile of the users.
4.1. Offline Experiments
In this subsection, we first present the problem formulation and the current model setup for the CTR prediction task on the search surface. Suppose that there is a potential product candidate eligible as ads when a user searches with a query string. The ranking model will take their (sparse, dense, cross) features as input and return a value between $[0,1]$ which is interpreted as the probability of the user clicking the ads (impression of ”sponsored product”). The model of choice in the current production environment is a Deep & Cross Network variant (Wang et al., 2017, 2021) trained on binary labels of historical user click engagements.
Next, we illustrate the integration of the projected embeddings of the product $embed_{p}$, query $embed_{q}$, and their derived features with the ranking model. Our model architecture as in Figure 3: a DCN tower that behaves the same as the current production model, and a new tower that takes the product and query embeddings from DashCLIP and their derived features as input. The intermediate outputs from the two towers are concatenated and passed through a few fully connected layers before the final sigmoid layer that produces the output prediction in the range from 0 to 1. In the offline experiment stage, we iterated over different architecture designs and found that this approach which separates the existing dense and sparse features from the new embeddings related features using the two-towers achieved the best performance. This architecture promotes the crossing between the different embeddings before interacting them with the existing features. The derived features are of two types: the similarity between embeddings and representations of user engagement history on the platform. The former is the cosine similarity of the embeddings such as $cosine(embed_{p},embed_{q})$ and aims to capture the present user intent. One particular instance of the latter is the $consumer\_p84d\_purchased\_product$ which is a list of product ids that the user purchased in the last 84 days on the platform. Then it retrieves the product embeddings using the product ids as indices from the pre-computed embedding table. A mean pooling averages the retrieved embeddings, and then pooled vector is fed to the model. Similarly, $cosine(embed_{p},pool(purchased\_product))$ captures how relevant the product is to the user’s purchase profile.

### v2

@@ 1. Introduction
Personalized recommendation (Cheng and Cantú-Paz, 2010; Pancha et al., 2022; Zhang et al., 2024; Baltescu et al., 2022), ubiquitous in e-commerce and social network platforms on the internet, is to find out relevant entities given the user preferences and context-dependent needs. This aligns with the business objective of creating user value and fostering long-term customer loyalty. Deep learning models have emerged as the foundation of modern machine learning systems in the space, excelling at capturing complex relationships between user preferences, product attributes, and contextual signals (Cheng et al., 2016; Guo et al., 2017; Wang et al., 2017, 2021). A key strength of deep learning models is their ability to learn embedding representations for categorical features while also leveraging pre-trained embeddings generated by more sophisticated models that are infeasible to serve online at web scale. In practice, deploying embedding-based approaches across various e-commerce applications presents several challenges: (1) Embeddings generated from historical user interactions may fail to capture the rich semantic structure of product attributes, resulting in biased exposure and diminished diversity. (2) Pretrained embeddings may struggle to generalize effectively to out-of-distribution scenarios, restricting their applicability across diverse product and user tasks.
@@
In this work, we present an embedding generation and alignment framework that we develop within the context of DoorDash, a large-scale online e-commerce platform which encompasses the delivery of groceries, retail products, alcohol, electronics, pharmaceuticals, and more. Our framework, named as DashCLIP, is built on top of the recent methodology of pre-training multi-modal models using contrastive learning (Radford et al., 2021; Li et al., 2022). DashCLIP leverages these key design choices to create semantic entity representations which are generalizable to different functional requirements:
@@
Multi-Modality Encodings: In e-commerce, products are typically associated with both textual and visual information. Instead of processing these modalities separately, it is essential to integrate the rich, yet unstructured, data from both sources to create a unified representation. In this work, we leverage contrastive learning on the product catalog of DoorDash to approximate human-like understanding of products, capturing the complementary information from each modality.
@@
Domain Adaptation: Transfer learning provides a solution by enabling fine-tuning of the off-the-shelf models for domain-specific applications, as demonstrated in prior work on fine-tuned image embeddings (Zhai et al., 2019) and fine-tuned text embeddings (Zhuang et al., 2020). To achieve this, we perform continual pre-training on a multi-modal transformer-based model using contrastive learning on our internal catalog data, leveraging product images and titles as inputs to adapt to the platform’s specific catalog distribution.
@@
Embedding Alignment: To ensure seamless integration with downstream applications, the architecture must be adaptable to accommodate new task objectives. For search recommendation, we introduce a second stage of alignment for a dedicated query encoder. It is designed to generate query embeddings that are learned in the same space as product embeddings, enabling more effective retrieval and ranking.
@@
Relevance Dataset Curation: For a product search application, product-query relevance is crucial to match the user intents and product offerings. Most prior systems resolve this connection through historical user engagements like clicks or conversions, which are prone to position and selection bias. To address this, we curate a high-quality relevance dataset that combines internal human annotations with knowledge from large language models, providing robust supervision for embedding alignment.
@@
The remainder of this paper is structured as follows: Section 2 covers the data collection and curation; Section 3 details the architecture of multi-modal encoders, along with the training procedure; Section 4 offers a thorough evaluation in search recommendation; Section 5 discusses potential e-commerce applications; and Section 6 & 7 concludes and reviews related work.
@@ 2.1. Catalog Dataset
The product catalog contains various types of information related to products available on DoorDash. This includes merchant-provided and internally collected data about the product. We curate a list of around 400k products and use their catalog data for our continual pre-training and evaluation tasks. While the catalog contains different fields, we find the following to be important due to their semantic value: 1) Title: The product name as shown to the user. 2) Image: A standardized product image. 3) Description: Usually, a single sentence elaborating some additional information, e.g. dimensions for packaged goods. 4) Detail: Usually, a few lines or a paragraph explaining the product features which are more verbose compared to the description. 5) Aisle Category: The product category classified with internal taxonomy, e.g. drinks, snacks.
For this paper, we only utilize the product title and image to train the embeddings. We make this decision for two reasons. First, the title and image are likely to receive the largest user attention on the platform as only these two fields are visible to the users in the category or search results page. Second, while the description and detail sometimes contain additional information, we observed some data quality issues with these two fields. The detail is often framed like an advertisement for the product which makes it quite noisy to use without further processing. Additionally, from an exploratory data analysis conducted on selected 13k products, we found that the description and detail to be missing for 32% and 20% of the products respectively, which raised data coverage concerns. We also use aisle category as part of the sampling strategy for contrastive learning and the additional downstream task described in Section 5.1.
@@ 2.2. Query Product Relevance Dataset
In order to align the query embedding and product embedding in the same space, we require a relevance dataset that assigns a relevance label from {0: irrelevant, 1: moderately relevant, 2: highly relevant} to each ¡query, product¿ pair. We refer to these labels as IR, MR, and HR. We used a hybrid approach to utilize both human annotation and a LLM to create such a dataset. Specifically, we have 700k human-annotated relevance labels, which is used to fine-tune a language model (GPT 3.5 as the final choice) to grade the relevance of any new ¡query, product¿ pair. In total, there are 32 million LLM-generated relevance labels for such pairs which cover more than 99% of our search volume. The distribution of the labels is {IR: 69%, MR: 20.5%, HR: 9.9%}. This process was done in late 2023 so GPT3, 3.5, 4 were evaluated. For our specific relevance task, we found that more sophisticated models may not guarantee enhanced performance, but fine-tuning has a notable improvement. Thus, we fine-tuned with 600k human annotations and evaluated it with the remaining 100k to select the best version for inference. The fine-tuning prompt specified the role of the language model and the definition of different relevance labels. After the initial prompt, a list of ¡query, product¿ examples with labels are also appended.
@@ 3.1. Model Architecture
The goal of DashCLIP is to create generalizable product representations that can be applied across various downstream applications. To achieve this, we adopt the vision-language pre-training framework outlined in (Li et al., 2022), which consists of two uni-modal encoders (one for image and one for text) and a third image-grounded text encoder. These encoders work together to generate product representations. To address scenarios involving user search, such as the search recommendation experiments discussed in Section 4, we incorporate an additional query encoder. This query encoder utilizes a text-only transformer to process normalized free-text search terms provided by the user. The final model architecture is illustrated in Figure 1.
@@ 3.2. Model Training
We initialize the image-text Product Encoders from a pre-trained checkpoint BLIP-14M (Li et al., 2022). The authors also make available the checkpoints which are fine-tuned on COCO (Lin et al., 2014) and Flickr30K (Young et al., 2014) datasets, however, we don’t opt for these as the fine-tuning datasets comprise of open-world and human images which are quite different from product images. After initialization, we adopt a two-staged training methodology:
@@ 3.2.1. Continual Pre-training of Product Encoders
In this first stage, we continue the pre-training of the Product Encoders on the 400k raw product images and titles from our Catalog Dataset mentioned in Section 2.1. By exposing the encoders to the product data from the platform’s catalog, the encoders will adapt to the characteristics and patterns of the product domain. This optimization stage aims to minimize the contrastive loss between image and text embeddings (ITC), and the matching loss of the image-text representations (ITM). We follow the same soft label creation and negative sampling strategies from (Li et al., 2022, 2021). In order to avoid overfitting the encoders, we freeze the first 8 layers of image and text encoders and early-stop at the lowest evaluation score of the matching loss.
@@ 3.2.2. Aligning Query and Product Encoders
where $C_{i}$ is the multi-modal hidden representation of the $i$-th product, $Q_{i}^{+}$ is the positive (relevant) query for the $i$-th product, $Q_{ij}^{-}$ is the $j$-th negative query among the $N$ negative samples for the $i$-th product. We average this loss over the batch size $B$. simsim is the cosine similarity function, and $\tau$ is the temperature parameter.
@@
3.2.3. Training cost
We perform these two training stages to obtain the final encoders as the inference model to generate DashCLIP embeddings for any product or query. The entire training was run on 16 Nvidia Tesla T4 (16 GB) GPUs. The total training time for one experiment run was roughly 10 hours. We are using AWS EC2 instances and the estimated cost for a single training run was $\approx$150 USD. To obtain the relevance labels, inference was done on OpenAI’s GPT 3.5, with an estimated cost of around 42k USD. Note that these labels are also being used for other applications within DoorDash’s ads funnel.
4. Search Ads Recommendation Experiments
The DashCLIP product and query embeddings are generalized for different downstream applications. In this section, we focus on the task of ads recommendation, which aims to find the most relevant advertisement for a user engaging the search surface of the platform. We are going to investigate how DashCLIP performs in the retrieval and ranking stages of the ads recommendation funnel, with comparisons with various common embedding approaches and ablation studies to investigate the best model architecture to make use of such embeddings. At the end, we deployed the best ranking candidate in the production environment to serve real users and verify the business impact. Highlights of the evaluations are as follow:
@@ 4. Search Ads Recommendation Experiments
    Off-the-shell vision-and-language models are not able to retrieve relevant products precisely. Alignment between the product and query in DashCLIP encoders is necessary to understand the e-commerce context.
  2. (2)
    DashCLIP embeddings outperform those learned through a randomly initialized embedding lookup matrix during end-to-end model training in ranking experiments. This highlights that utilizing the semantic information of products and queries to pre-train the embeddings provides a strong prior for downstream applications.
@@
4.1. Retrieval
In the retrieval task, we leverage the embedding of a user’s query to perform a K-nearest neighbor search in the embedding space of the product to create a ranked list of potential relevant candidates for the next downstream selection (ranking etc). Both effective representations and alignment between them are essential to retrieve relevant items at top positions. For subsequent evaluation, we used the hold-out subset of Query Product Relevance Dataset, which consists of ¡query, product¿ pairs that are not used in the training of DashCLIP, as the seed queries and corresponding retrieval pool of product candidates.
Methods to compare: For CLIP (Radford et al., 2021), we used the ViT-B/32 Transformer architecture as the image encoder for product image and the masked self-attention Transformer as the text encoder for both user query and product title. The encoders are initialized from Hugging Face’s checkpoint. For BLIP, we used the unimodal image and text encoders in the same way after the final projection layers, thus representations are aligned in the same space by the contrastive loss. Note that in BLIP, the representation from the image-grounded text encoder is not aligned with the unimodal encoders. We used the pretrained BLIP-14M checkpoint (Li et al., 2022). For FLAVA (Singh et al., 2022), the setup is similar to BLIP and we used the checkpoint from Hugging Face. For DashCLIP, the natural choice is the embedding from the image-text encoder and query encoder which is aligned by Product-Query contrastive (PQC) loss.
From Table 1, DashCLIP outperformed all baselines by significant gains, demonstrating the effectiveness of product-query alignment in our proposed framework. General LLM, while much better than random with the power of world knowledge, lacks the specificity of the e-commerce domain and often fails for short but specific queries. Surprisingly, CLIP (Image) performs notably well in terms of precision. We found that for many branded queries (such as ‘celsius drink’), the image encoder is able to recognize the text on the product image to perform keyword matching as a zero-shot OCR task. We also experimented with multi-modal embeddings of BLIP and FLAVA. Since there is no alignment between this representation and the text-only encoding of the query, the performance is significantly worse by default.
Embedding Variant Graded Rel Recall@K Precision@K
Random 0.102 0.194 0.027
CLIP (Text) 0.346 0.530 0.164
CLIP (Image) 0.529 0.666 0.325
BLIP (Text) 0.328 0.457 0.190
BLIP (Image) 0.316 0.468 0.178
FLAVA (Text) 0.276 0.446 0.137
FLAVA (Image) 0.393 0.562 0.230
DashCLIP (Image-Text) 0.667 0.743 0.515
Table 1. Search Retrieval Experiment. Evaluation is done on the top-10 retrieved product candidates. Graded Relevance considers the three-classed relevance label {IR, MR, HR} while Recall@K and Precision@K treats MR+HR as the binary positive label and IR as the negative label.
@@ 4.1. Retrieval
NWNW $\setminus$ $U_{PurcHist}$ DCN baseline .7456 $\pm$ .0004 .9183 $\pm$ .0006
@@
Table 2. Search Ranking Experiment. We collect the evaluation data from one week after the training data (called NWNW). Several evaluations were done to highlight different perspectives of the performance gain. The best candidate with product + query + purchase history embeddings is bold and its improvements over the baseline were statistically significant ($p<0.05$). Also, the candidate performed better for users with purchase history ($U_{PurcHist}$) by capturing a semantic profile of the users.
4.2. Ranking: Offline Experiments
In the ranking task, we first present the problem formulation and the current model setup for the CTR prediction task on the search surface. Suppose that there is a potential product candidate eligible as ads when a user searches with a query string. The ranking model will take their (sparse, dense, cross) features as input and return a value between $[0,1]$ which is interpreted as the probability of the user clicking the ads (impression of ”sponsored product”). The model of choice in the current production environment is a Deep & Cross Network variant (Wang et al., 2017, 2021) trained on binary labels of historical user click engagements.
Next, we illustrate the integration of the projected embeddings of the product $embed_{p}$, query $embed_{q}$, and their derived features with the ranking model. Our model architecture is shown in Figure 3: a DCN tower that behaves the same as the current production model, and a new tower that takes the product and query embeddings from DashCLIP and their derived features as input. The intermediate outputs from the two towers are concatenated and passed through a few fully connected layers before the final sigmoid layer that produces the output prediction in the range from 0 to 1. In the offline experiment stage, we iterated over different architecture designs and found that this approach, which separates the existing dense and sparse features from the new embedding-related features using the two-towers, achieved the best performance. This architecture promotes the crossing between the different embeddings before interacting them with the existing features. The derived features are of two types: the similarity between embeddings and representations of user engagement history on the platform. The former is the cosine similarity of the embeddings such as $cosine(embed_{p},embed_{q})$ and aims to capture the present user intent. One particular instance of the latter is the consumer_p84d_purchased_productconsumer\_p84d\_purchased\_product which is a list of product ids that the user purchased in the last 84 days on the platform. Then it retrieves the product embeddings using the product ids as indices from the pre-computed embedding table. A mean pooling averages the retrieved embeddings, and then pooled vector is fed to the model. Similarly, $cosine(embed_{p},pool(purchased\_product))$ captures how relevant the product is to the user’s purchase profile.